impure copper obtained from minerals contains impurities as iron, gold and silver. after passage of 200 a current for 400s, the mass of anode decreased by 27g and that of cathode increased by 24.5g. calculate the respective percentage of iron and copper originally present (given weight of cu = 63.5g and fe = 56g)


1
Expert's answer
2021-07-26T06:32:24-0400

Molar Mass of Gold = 196.97

Molar Mass of silver = 107.87

63.5 - 27 = 36.5

56-24.5 = 31.5

372.84

Iron = 56/372.84 × 100 = 15.02%

Copper = 63.5/372.84 × 100 = 17.03%